OpenStudy (anonymous):

Yasmeen is setting up a business. She borrows $5000 from a loan company. The loan company charges 6% per year simple interest. How much interest will Yasmeen pay after 3 years?

OpenStudy (anonymous):

so do you know how much after one year?

OpenStudy (anonymous):

NOPE

OpenStudy (anonymous):

what is 6% or 5000?

OpenStudy (anonymous):

*of

OpenStudy (anonymous):

what is 6% of 5000?

OpenStudy (anonymous):

300

OpenStudy (anonymous):

good so that is added on after one year

OpenStudy (anonymous):

so how much is it after one year?

OpenStudy (anonymous):

your genies,just tell me the answer :)

OpenStudy (anonymous):

you have to learn

OpenStudy (anonymous):

its easy

OpenStudy (anonymous):

5000 + 300 =

OpenStudy (anonymous):

the answer is 900

OpenStudy (anonymous):

cluo your so nice, just tell me the answer :)

OpenStudy (anonymous):

.....?

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

its (5000 + 300) * 6% + (5000 + 300) = x (x * 6%) + x x - 5000 = answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Can i ask more questions?

OpenStudy (anonymous):

what?